Solve



x^2y′′+ xy′− y =1/x + 1

1
Expert's answer
2022-03-16T12:52:18-0400

"a_0x^ny^n+a_1x^{n-1}y^n+...+a_{n-1}xy'+a^ny=f(x)"

"x=e^u"

"a_0\lambda(\lambda-1)(\lambda-2)...(\lambda-n+1)...+a_{n-2}\lambda(\lambda-1)+a_{n-1}\lambda+a_n=0"

"(\lambda-1)\lambda+\lambda-1=0"

"\lambda^2-1=0"

"y''-y=\frac{1}{e^u}+1"

"\lambda_1=1"

"\lambda_2=-1"

"y=Ce^u+\frac{C_1}{e^u}"


"y_i=u^se^{au}(R_m(u)cos(bu)+T_m(u)sin(bu))"

s=0 if a+bi is not a root or s=k if it is.

Solution for 1:

a+bi=0, then s=0

"y_0=A"

"y''_0=0"

A=-1

"y_0=-1"

Solution for "\frac{1}{e^u}"

a+bi=-1

s=1

"y_1=\frac{Au}{e^u}"

"y''_1=\frac{Au-2A}{e^u}"

"\frac{-2A}{e^u}=\frac{1}{e^u}"

A=-0.5

"y_1=-\frac{u}{2e^u}"

"y=Ce^u-\frac{u}{2e^u}+\frac{C_1}{e^u}-1"

u=ln(x)

"y=\frac{-ln(x)}{2x}+Cx+\frac{C_1}{x}-1"